OpenStudy (praetorian.10):

The equation v = u + at, is used to calculate velocity. Calculate the velocity (v) if u = 2.6, a = 1.2 and t = 5 6.6 13.85 8.6 69

OpenStudy (praetorian.10):

@texaschic101 @sujuexob2st

OpenStudy (praetorian.10):

8.6?

OpenStudy (praetorian.10):

i think :/

OpenStudy (anonymous):

v=8.6

OpenStudy (praetorian.10):

yay i was right :D

OpenStudy (texaschic101):

v = u + at u = 2.6, a = 1.2, t = 5 now just sub v = 2.6 + (1.2)(5) v = 2.6 + 6 v = 8.6

OpenStudy (praetorian.10):

CHEERS :D